OpenStudy (anonymous):

Say that a and b are the lengths of the leg of a right triangle and c is the lenght of the hypotenuse. find c if a = 15 and b=20

OpenStudy (solomonzelman):

\[a^2+b^2=c^2\]where "a" and "b" are the legs of the triangle and "c" is the hypotenuse

OpenStudy (solomonzelman):

you can write this formula for c, as: \[\sqrt{a^2+b^2}=c\]

OpenStudy (anonymous):